### v1

# Probability calibration for precipitation nowcasting

Abstract
Reliable precipitation nowcasting is critical for weather-sensitive decision-making, yet neural weather models (NWMs) can produce poorly calibrated probabilistic forecasts. Standard calibration metrics such as the expected calibration error (ECE) fail to capture miscalibration across precipitation thresholds. We introduce the expected thresholded calibration error (ETCE), a new metric that better captures miscalibration in ordered classes like precipitation amounts. We extend post-processing techniques from computer vision to the forecasting domain. Our results show that selective scaling with lead time conditioning reduces model miscalibration without reducing the forecast quality.
1 Introduction
Precipitation nowcasting—forecasting the immediate future with lead times of up to four hours at high temporal and spatial resolution— supports time-sensitive decisions in disaster response, transportation safety, urban drainage management, and winter road maintenance gen-models ; nowcast-challenge ; nowcastnet . As climate change drives shifts in precipitation patterns and increases the frequency of extreme weather events having skillful nowcasts is ever more important.
Recent years have seen rapid advances in precipitation forecasting through deep neural networks (DNNs) shi2025deeplearningfoundationmodels ; aabonev2023modelling ; mcnally2024datadrivenweatherforecasts ; bi2023accurate . Neural weather models (NWMs) are state-of-the-art systems and are being deployed operationally by both industry and meteorological agencies andrychowicz2023deeplearningdayforecasts ; lang2024aifsecmwfsdatadriven ; adrian2025data . Many applications demand not only accurate but also probabilistic forecasts, where predicted probabilities reflect the true likelihood of events. Two main approaches exist: generating ensembles of perturbed predictions, or directly building probabilistic models bi2023accurate ; andrychowicz2023deeplearningdayforecasts ; price2025probabilistic . In this work, we focus on the latter.
A key requirement for reliable probabilistic forecasts is calibration—the alignment between predicted probabilities and observed event frequencies. For a model that predicts a class $\hat{Y}$ with probability $\hat{P}$, perfect calibration is formally defined as $\mathbb{P}(\hat{Y}=Y|\hat{P}=p)=p,\ \text{for all}\ p\in[0,1]$ and all class labels $Y\in\left\{0,\dots,K-1\right\}$ Nixon_2019_CVPR_Workshops . For classification and semantic segmentation models, calibration is often assessed via the expected calibration error (ECE)
$\operatorname{ECE}=\sum_{b=1}^{B}\frac{n_{b}}{N}\left|\operatorname{acc}(b)-\operatorname{conf}(b)\right|,$ (1)
where $n_{b}$ is the number of predictions in bin $b$, NN is the total number of data points, and $\operatorname{acc}(b)$ and $\operatorname{conf}(b)$ are the mean accuracy and confidence in that bin Nixon_2019_CVPR_Workshops ; guo ; wang10205155 ; NEURIPS2019_f8c0c968 . However, ECE has well-known shortcomings, especially in multiclass problems with ordered categories, such as precipitation amounts Nixon_2019_CVPR_Workshops ; gawlikowski2023survey . In this context, the metric’s focus on the predicted class and its associated confidence obscures miscalibration across the full range of precipitation bins. For example, a winter maintenance operator must understand probabilities across multiple thresholds (e.g., 1 mm versus 10 mm of snowfall), not just the most likely category. Static calibration error (SCE) Nixon_2019_CVPR_Workshops extends ECE to a multiclass setting but it’s still intended only for independent classes. Therefore, for a model predicting a probability vector $\hat{P}(r)$ over precipitation rates $r\in[R_{1},\dots,R_{K}]$, the calibration is better defined by
$\mathbb{P}(r>R\ |\ \hat{P}(r>R)=p)=p,\ \text{for all}\ p\in[0,1],\ R\in[R_{1},\dots,R_{K}].$ (2)
In other words, given 100 predictions for precipitation $r>R$ each at confidence 0.8, we expect that for 80 % of those predictions, precipitation will exceed $R$.
In this note, we address the miscalibration of probabilistic NWMs in precipitation nowcasting. We introduce the expected thresholded calibration error (ETCE) as a more appropriate metric for assessing miscalibration in our setting. We study post-processing techniques designed to adjust model confidence values to better match observed precipitation frequencies. While there is active research on the topic, existing calibration methods are limited—particularly in computer vision—and, to our knowledge, absent in the context of NWMs 9710154 ; wang10205155 ; ding_local .
2 Methodology
2.1 Expected thresholded calibration error
To estimate the calibration error of Eq. 2, we first compute confidences for the thresholded prediction $\hat{P}(r>R_{k})$ for all $k\in[1,\dots,K]$. Then, for each threshold we bin the predictions by predicted confidence into $B$ evenly spaced bins. Finally, for each bin $b$ and threshold $R_{k}$ we compute the mean predicted confidence $\operatorname{conf}(b,R_{k})$ and mean accuracy $\operatorname{acc}(b,R_{k})$. Then, we compute the difference between predicted confidence and accuracy as
$\operatorname{ETCE}=\frac{1}{K}\sum_{k=1}^{K}\sum_{b=1}^{B}w_{b}\left|\operatorname{acc}(b,R_{k})-\operatorname{conf}(b,R_{k})\right|.$ (3)
where $w_{b}$ are bin weights. In this application, we apply uniform weighting $w_{b}=1/B$ as precipitation is a rare event and weighting by the number of samples would emphasize dry events in the metric. In our experiments we have enough data to select $B=20$ (See Figure A.1 in the Appendix).
Figure 1: Miscalibration diagram at precipitation threshold $1.5\operatorname{mm/h}$ and selected confidence bins [0.05,0.10],[0.45,0.50],[0.85,0.90][0.05,0.10],\ [0.45,0.50],\ [0.85,0.90]. The mean confidence and mean observed frequency for each bin are depicted by dashed and solid curves respectively.
To build intuition on ETCE, we illustrate the miscalibration between predicted confidence and observed frequency at fixed threshold $R_{k}=1.5$ mm/h for three selected bins in Figure 1. The calibration error $|\operatorname{acc}(b)-\operatorname{conf}(b)|$ corresponds to the filled area between observed frequency and confidence. In this example, the observed frequency is smaller than the average confidence in the same bin meaning that the model is overconfident. For a better calibrated model, the area between observed frequency and predicted confidence is smaller.
2.2 Calibration methods
In the literature there are multiple post-processing tools for calibrating probabilistic models in deep learning. Here we describe the calibration methods we extended and tested in the forecasting domain.
Temperature scaling (TS), and its variations, is calibration method which has been shown to be effective for classification tasks guo ; hinton2015distillingknowledgeneuralnetwork ; balanya2024adaptive . In TS, a single parameter $T\in\mathbb{R}^{+}$ is learned—typically by minimizing the negative log-likelihood—to scale the predicted probability $\boldsymbol{\hat{p}}=\sigma_{\operatorname{softmax}}(\boldsymbol{z}/T)$. In segmentation tasks, one temperature is optimized to scale predicted probabilities of pixels and all samples.
Local temperature scaling (LTS) ding_local is an extension of TS in which a different temperature is applied for each pixel $x$ in a sample. In this approach, we need to learn a regressor for mapping a logit vector $\boldsymbol{z}$ to a temperature value $T$. LTS is proposed to better calibrate segmentation models where especially the label boundaries are often miscalibrated. In this work and in ding_local , LTS learns the temperature mapping using a small hierarchical CNN. Different from ding_local , we only use the predicted logits as input and also condition the regressor with lead time using Feature-wise Linear Modulation (FiLM) perez2017filmvisualreasoninggeneral which applies affine transformations to intermediate feature maps based on external information.
Selective scaling (SS) wang10205155 is based on the observation that the major cause of neural network miscalibration is overconfidence on mispredictions. SS uses a classifier on the logits $\boldsymbol{z}$ to first detect mispredictions of the base model, and then applies scaling with temperature $T>1$ only on the mispredictions to reduce overconfidence to obtain the calibrated probability vector $\boldsymbol{\hat{p}}$,
[…]
In wang10205155 the classifier is a 3-layer MLP conditioned on the logits. We augment that architecture by using FiLM to pass the lead time encoding. We also investigate enhancing the classifier’s spatial view through larger attention-based architectures.
2.3 Data and model
Base model. In our experiments we hold the probabilistic base model fixed. The model is conditioned on lead time so the predictions are independent across lead times. The architecture has three main components: a spatial encoder consisting of a sequence of convolutional layers to downsample the input data from $512\times 512$ pixels to $64\times 64$ with 512 channels; an attention block composed of four axial-attention layers axial outputting 512 channels; and finally a classification head outputting 12 channels, one for each precipitation rate bin. The total number of weights is 21M. The base model output logits are used as input for training the calibrator models.
Data. The input data consists of 7 steps of MRMS radar images, 2 steps with 16 channels of GOES satellite, 1 step of precipitation prediction by HRRR (numerical weather model in North America); as well as topography, longintude, latitude, temporal information, and lead time. The target is MRMS discretized to 12 bins from 0.2 mm/h up to 10+ mm/h and one-hot encoded such that the base model predicts a 12-vector as a probability over the binned precipitation rates.
Calibration training data. The data used in the development and evaluation of the calibration methods is temporally non-overlapping from the data used for training of the base model. We use 110K unique samples (each sample has spatial size $64\times 64$) to train the classifiers for flagging mispredictions in selective scaling. To optimize the temperatures in temperature scaling, selective scaling, and for learning the logit-temperature mapping in local temperature scaling, 1K samples are used. Finally, the uncalibrated and calibrated ETCE scores are computed over 47K samples not included in the training data of the calibrators.
3 Results
We summarize the ETCE scores averaged over the lead times in Table 1. The first row corresponds to the base model. We include the number of learnable weights for the calibration methods; it includes classifier weights for the selective scaling schemes. The MLP-based selective scaling calibrator yielded the best improvement to model calibration with 23% ETCE reduction. The Segformer-based selective scaling calibrator came as a close second with a 21% reduction. Although temperature scaling has shown positive results in some computer vision problems, it did not reduce miscalibration here. Local temperature scaling was detrimental to model calibration, but the damage it did to calibration was alleviated when lead time encoding was used.
Calibrator num. params F1-score avg. ETCE $\Delta$ ETCE (%)
Uncalibrated - 0.565 0.079 -
Temperature scaling 1 0.565 0.080 -1.0
LTS (no lead time cond.) 2,107 0.573 0.096 -21.3
LTS 2,143 0.564 0.082 -3.6
Selective scaling w/ MLP 3,254 0.564 0.060 23.5
Selective scaling w/ Segformer B0 3,728,550 0.567 0.062 21.6
Table 1: ETCE scores averaged over all lead times for different calibrator models and the relative improvement over the uncalibrated baseline model. We also show the average F1-score computed for thresholded precipitation at 1 mm/h.
A more detailed ETCE comparison per lead time is shown in Figure 2. Selective scaling reduced miscalibration effectively over all lead times. At shorter lead times up to 150 minutes, selective scaling with either MLP or Segformer-B0 classifiers reduced ETCE equally. But at longer lead times the MLP-based calibrator performed best.
Figure 2: ETCE as a function of lead time for the uncalibrated model, and after applying temperature scaling, local temperature scaling and selective scaling.
We also tested more complex classifiers with the selective scaling approach—Segformer-B1 and B2—and found only minor improvement in ETCE compared to MLP and B0 classifiers (See Figure A.2 in the Appendix). The marginally increased accuracy with a Segformer-B2 classifier does not justify the massive increase in model complexity. A more detailed look on miscalibration reduction is shown in Figure A.3 in the Appendix.
4 Conclusions
We introduced the expected thresholded calibration error (ETCE) to measure probability calibration of probabilistic models in forecasting. Different from standard computer vision tasks, in forecasting there is the lead time dimension. By combining selective scaling with lead time encoding we reduced the base model’s calibration error by up to 23.5 %. Based on these results, future works should build on selective scaling, possibly by conditioning the calibrator on spatial and/or temporal information in addition to lead time.
Appendix A Appendix
A.1 Number of predictions in different confidence bins
Figure A.1: Number of predictions at threshold $r\geq 1.0$ mm/h within five selected confidence bins.
Figure A.1 shows the number of predictions in five selected confidence bins. This example is for thresholded predictions at $r\geq 1.0$ mm/h.
A.2 Selective scaling with Segformer classifiers
Figure A.2: ETCE as a function of lead time for selective scaling using different classifiers for flagging mispredictions. Uncalibrated baseline shown with a dashed line.
Figure A.2 shows ETCE as a function of lead time for selective scaling with different classifier models compared against the base model. Overall, using Segformer-B2 as the classifier resulted in approx. 1.3 % lower ETCE compared to using the MLP but the significantly larger computational cost of the Segformer makes the MLP a more viable choice.
A.3 Calibration effect on ETCE
Figure A.3: The difference between model confidence and accuracy for thresholds $r\geq 1.0$ mm/h and $r\geq 2.0$ mm/h before and after applying calibration (Selective scaling with MLP classifier). Reduction in colored area shows reduction in miscalibration. Smallest area in between dashed and solid lines is best.
Further detail on ETCE improvement under calibration is shown in Figure A.3. Here, the specific calibration is selective scaling which we use with the MLP classifier for flagging mispredictions. The colored areas—gray for uncalibrated and green for calibrated—show the difference between average model confidence and accuracy at three thresholded confidence bins. For the uncalibrated baseline, we observe overconfidence across lead times and precipitation thresholds. The figure also shows that selective scaling reduces miscalibration effectively which shown by the smaller colored area. This is especially true when the predicted confidence is high.